The total number of electrons in 1.6g of CH4 to that in 1.8 g H2O is

Grade:12th pass

1 Answers

Antara
3 years ago
No of moles of CH4=1.6/16=0.1
No of moles of H2O=1.8/18=0.1
no of electrons in 1 mole of CH4=10 (6 e of C and 1 e of each H)
no of electrons in 1 mole of H2O=10 (8 e of O and 1 e of each H)
So total no of e’s in 1.6g CH4=10*no of moles=10*0.1=1
total no of e’s in 1.8g  H2O=10*no of moles=10*0.1=1
So the required ratio=1:1
 
Hope it helps :)